OpenStudy (anonymous):

Solve the system of equations: y= x + 3 y= 2x + 4

OpenStudy (eyust707):

there are two different methods you could use

OpenStudy (anonymous):

the substitution method

OpenStudy (eyust707):

ok

OpenStudy (eyust707):

so up there you clearly wrote that y is equal to "x + 3" right?

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

so couldnt we go to the other one and change y to "x+3"

OpenStudy (anonymous):

oh ok yes we can

OpenStudy (eyust707):

so x+3 = 2x + 4 I just substituted "x+3" in for y =P

OpenStudy (eyust707):

solve for x

OpenStudy (anonymous):

so it would be -1 for x

hero (hero):

I showed him the best method to use and he still didn't get it

OpenStudy (eyust707):

yep so x= -1

OpenStudy (anonymous):

i understand this method hero

OpenStudy (eyust707):

now they usually also want y

OpenStudy (anonymous):

yes so i will solve for y right now

OpenStudy (anonymous):

would you plug in -1 for the y because for the method that our teacher taught us we plugged in the number that we got for x and used it in y

OpenStudy (anonymous):

o no nevrmind

OpenStudy (eyust707):

x= -1 no trickery =)

OpenStudy (anonymous):

y = 4

OpenStudy (eyust707):

y = x + 3

OpenStudy (eyust707):

y = (-1) + 3

OpenStudy (eyust707):

y =???

OpenStudy (anonymous):

y=-1+3

OpenStudy (anonymous):

2

OpenStudy (eyust707):

ding ding ding

OpenStudy (anonymous):

oh ok now im starting to understand it.

OpenStudy (eyust707):

thats the goal!

OpenStudy (anonymous):

yes thank you so much

OpenStudy (eyust707):

what we have just done is found an x and found a y that satisfies both those equations

OpenStudy (anonymous):

yes so that is easier than graphing in my opinion

OpenStudy (eyust707):

yep way easier.. butt!! if you did graph it and we went to the point (-1, 2) what special thing would be happening?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

let me think a moment

OpenStudy (eyust707):

so above those equations represent two lines

OpenStudy (anonymous):

yeah

hero (hero):

You could easily understand my method. And I showed you the steps so you know they are simple.

OpenStudy (anonymous):

i obviously got the answer with eyust's method and not for your method.

OpenStudy (eyust707):

we found the point when both lines are equal

OpenStudy (anonymous):

oh ok

OpenStudy (eyust707):



OpenStudy (eyust707):

we found that point!

OpenStudy (eyust707):

right when they are equal!

OpenStudy (anonymous):

yeah we did

OpenStudy (eyust707):

we call this the point of intersection

hero (hero):

Which is exactly why when I solve them I write the equations in terms of y first

OpenStudy (anonymous):

oh ok eyust

OpenStudy (eyust707):

everyone learns math differently... seems like yugo understood this method easier

hero (hero):

y= x + 3 y= 2x + 4 Do you realize you can graph both equations on the same graph using a calculator to check your answer?

hero (hero):

eyust, let me have a shot please. You've spoke with him all this time.

OpenStudy (anonymous):

eyust showed me a process where i understand better than yours. There is nothing wrong with yours either but you have to realize that some people understand different things hero.

hero (hero):

I know you want to gloat because you helped him learn your method

OpenStudy (eyust707):

lol see ya around guys

OpenStudy (anonymous):

haha ok thanks eyust

hero (hero):

Yugo, nothing is preventing you from learning my method. If you just understand that when you set both equations equal to y y = y for example, you are finding the point where both lines are equal. I know it may seem like my method is difficult to understand, but actually, it is much simpler, and also, the other methods don't always work. You want something that is consistent and my method is much more consistent and reliable.

OpenStudy (anonymous):

Thats true and makes since. Can I try an example of your method or use a previous problem